OpenStudy (anonymous):

How do you solve the rational equation 1/2 + 1/4 = 1/x?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (kaylaprincess):

1/2 + 1/4 = 3/4 = 1/(4/3) x = 4/3

OpenStudy (kaylaprincess):

So you have: (3/4) = 1/x. This means that the number x has as its reciprocal (3/4). So to find the number, take the reciprocal of the reciprocal. This is then : x = 4/3